# Improving Concept Alignment in Vision-Language Concept Bottleneck Models

### v1

@@ 4.2 Improving Concept Alignment
Hence, to improve the faithfulness and reliability of VL-CBM, we apply the concept projection later (bootstrap) and optimize the model with our CSS objective (CSS VL-CBM). Our CSS method not only substantially increases the concept accuracy, it also closes the gap with black-box methods on image classification performance. Especially for the CUB dataset, our CSS method surpasses the black-box classifier. The “number of concept labels per class” used to train the CSS VL-CBM model for the above datasets are as CUB - 9 labels (30%), RIVAL - 8 labels (0.1%), and AwA2 - 10 labels (1.4%). The values within brackets denote the “concept labels used for training as a percentage of the total concept labels provided by the dataset”. The results show that even with such small percentages of concept labels, our CSS method can substantially increase the concept accuracy. The enhanced and faithful concept scores after improving concept alignment with CSS is presented in Fig. 2 (CSS Concept Scores).
@@
WBC Attribute prediction. WBCAtt contains White Blood Cell (WBC) images annotated with 31 “cell morphology” attributes that aids in diagnosis in Hematology. We directly evaluate the models on their attribute prediction accuracy and do not perform any image classification. While contemporary approaches [48, 46] use the HAM10000 [35] dataset to study VL-CBM in the medical domain, the availability of attribute labels in WBCAtt makes the latter suitable for our work. For our evaluation, we train the PLIP model [9] (a VLM model for Pathology AI), with 50% of the attribute labels. From the table, we can see that our CSS method substantially increases the attribute prediction accuracy and closes the gap with black-box attribute predictor.
Figure 3: Number of concept labels vs CSS VL-CBM performance.
Impact of the number of concept labels. Since our approach is semi-supervised, we further investigate the impact of the number of human-annotated concept labels on the CSS VL-CBM performance in Fig. 3.
For CUB and RIVAL datasets we progressively increase the concept labels used for training, for which we observe a proportional increase in the classification and concept accuracies.
For WBCAtt, we investigate how the black-box (PLIP vision encoder) attribute predictor compares against the CSS (PLIP) VL-CBM attribute predictor, for an increase in the attribute labels used for training. Note that there is no image classification task for WBCAtt. We can see that our CSS method outperforms black-box attribute predictor in a low data-resource setting.
4.3 Comparison with other methods
CUB. The CUB dataset is extensively used by the Explainable AI community to build and study concept-based interpretable models. Hence, we comprehensively compare our Contrastive Semi-Supervised VL-CBM method with contemporary VLM-based interpretable models on the CUB dataset. The results are presented in Table 3. For a fair comparison with LF-CBM [25] and PCBM [48], we perform two-stage training. In stage 1, a ResNet18 backbone with a linear concept predictor is trained with the concept labels generated by our CSS VL-CBM model. It minimizes the $L_{concept}$ loss function. In stage 2, a sparse-linear SGD classifier from scikit-learn is trained with 5000 steps as described in [48]. The backbone, concept predictor, and sparse-linear classifier together constitute a CBM.
Table 3: Comparison of CSS VL-CBM with contemporary methods on CUB.
Method Base CBM Training Classifier Class (%) Concept (%) LF-CBM [25] ResNet-18 2-stage Sparse-Linear 74.31 - PCBM [48] ResNet-18 2-stage Sparse-Linear 59.6 - CLIP VL-CBM ResNet-18 2-stage Sparse-Linear 68.77 16.53 CSS VL-CBM (Ours) ResNet-18 2-stage Sparse-Linear 72.85 62.8 CLIP VL-CBM ResNet-18 joint Linear 74.91 18.84 CSS VL-CBM (Ours) ResNet-18 joint Linear 80.36 63.17 CompDL [49] CLIP ViT-B/32 - Linear 52.6 - LaBo (full shot) [46] CLIP ViT-B/32 - Linear 81.0 - CSS VL-CBM (Ours) CLIP ViT-B/16 joint Linear 81.45 63.53
From the table, we can see that the classification accuracy of ResNet-18 CBM supervised by our CSS VL-CBM approach is on par with LF-CBM and achieves much higher performance than PCBM. However, we observe the best performance with joint training of the backbone, concept predictor, and linear classifier layers with $L_{concept}$ and $L_{ce}$ losses. Grad-CAM [31] visualization of the concepts learned by the (jointly trained) ResNet18 CBM supervised by our CSS VL-CBM is presented in Fig. 4. From the figure, we can see that the improved concept knowledge (activating the truthful concepts for a given image) of our CSS VL-CBM model was effectively transferred to the ResNet model. It is able to correctly associate a color or texture concept to the corresponding bird part. Also, our method performs better than LaBo [46] with just $312$ (CUB) expert concepts, while LaBo utilizes $10,000$ LLM-generated bottleneck concepts.
Figure 4: Grad-CAM visualization of the learned concepts.
@@ 4.3 Comparison with other methods
Table 4: Comparison of CSS VL-CBM with other methods on RIVAL and AwA2.
Dataset RIVAL AwA2 CBM Method  Text2Concept   CSS  PCBM   ECBM   CSS Classification Accuracy (%) 95.38   98.86  88   91.2   93.2 Concept Accuracy (%) -   72.84  71.9   85.84   81.3
@@
RIVAL and AwA2. We further compare CSS VL-CBM with contemporary methods on the animal classification datasets in Table 4. For RIVAL dataset, CSS achieves a higher classification performance compared to Text2Concept [22]. For AwA2 dataset, CSS achieves a higher classification accuracy compared to other methods, but achieves a slightly lower concept accuracy compared to ECBM. It is important to note that PCBM [13] and ECBM [43] do not use “text” concepts and they optimize the feature extractor (vision backbone) during training.
5 Intervention
@@
When a CBM makes an error, one can inspect the incorrect predictions and manually tune the classifier weights or modify the concepts to fix those errors. This approach is called “intervention” or “model editing”. Conventional approaches to intervention are typically done at the instance-level [25, 48, 14], where a handful of error images are arbitrarily chosen for intervention. However, for fine-grain multi-class classification problems like CUB which have more classes (some bird species are visually similar) and relatively fewer samples per class, choosing the error images is non-trivial. Moreover, it is desired to perform intervention such that it increases the performance globally [48] instead of rectifying the errors for a few images. Hence, we propose a class-level intervention procedure where we first identify the confounding classes and then choose the error images of these classes for intervention. Confounding classes are bird species, that are visually similar yet semantically different and the model makes the most errors for these challenging classes. This class-level intervention is done in the following steps.
Figure 5: Confounding classes - birds that are visually similar yet belong to different sub-species.
@@
Before Interv.   After Interv. Total error for CG (30) 27 14 Total error for WG (30) 15 9 CG misclassified as WG 13 4 WG misclassified as CG 3 1 Total error for CT (30) 21 11 Total error for AT (29) 3 2 CT misclassified as AT 12 6 AT misclassified as CT 0 0 Overall Classification Accuracy (%) 81.45 82.57
Step 1: Identify the confounding classes. First, we compute a $K\times K$ error matrix by evaluating the trained CSS VL-CBM model on the test set, where $K$ is the number of classes. The error matrix gives the error distribution for every class and the index of the matrix element with the highest value (error) gives the confounding class pairs. In our experiments, we identified two pairs of confounding classes “California Gull - Western Gull” and “Common Tern - Artic Tern” as shown in Fig. 5. The total number of classification errors for these confounding classes before intervention are presented in Table 5.
Step 2: Inspecting the concept space. Next, we inspect the concept scores of the error images in these confounding class pairs. We find that the concept score distribution of the confounding classes are nearly identical, which makes it difficult to discriminate them. Our model suffers from Type-2 error [25] i.e., the number of bottleneck concepts and their score distributions are insufficient to effectively discriminate the samples.
Step 3: Expanding the concept set. To mitigate the Type-2 error and to improve the overall performance, we expand the concept set by leveraging the concepts provided by LaBo [46]. Note that the new concepts to be added should be visually groundable by the CLIP model and hence we use the “filtered” concept set of LaBo. We choose $top-32$ concepts per confounding class and expand the concept set from the original 312 concepts to 440 concepts. Formally, let $T^{\prime}$ be the new set of 128 text concepts and $C^{\prime}_{i}=E_{I}(x_{i})\cdot E_{T}(T^{\prime})^{\intercal}$ be the CLIP concept scores for it. Now the concept scores for all the 440 concepts for an image $x_{i}$ is given by $C^{\prime}=$ Concat($C_{i},C^{\prime}_{i}$). During this intervention procedure we do not train a new concept projection layer for the expanded concept set.
@@
6 Limitations
Despite the numerous advantages, VL-CBMs have two potential limitations.
Ineffable concepts. Though natural language offers a convenient way to build interpretable models with high-level abstract concepts, their expressive power can be limited i.e., the subtle visual cues needed required for solving certain classification tasks, like face recognition, might be difficult to express in words.
Unknown concepts. All CBMs (including VL-CBMs) assume that the complete set of salient concepts required for solving the classification task is known prior to training. However, this requirement cannot be satisfied for all use cases. In such scenarios, prototype-based methods with learnable visual feature prototypes are suitable where they can “discover” unbeknownst concepts during training.
7 Conclusion
In this paper, we investigated the faithfulness of VL-CBM for expert-defined concepts and found that VLMs like CLIP have poor concept alignment. Hence, we proposed a novel Contrastive Semi-Supervised approach that improved the concept alignment in VL-CBM with fewer human-annotated concept labels. We also introduced a class-level intervention procedure for fine-grain classification problems that reduced the error for the confounding classes to increase the overall performance of the model.
8 Acknowledgement
This work was carried out at the Rapid-Rich Object Search (ROSE) Lab, School of EEE, Nanyang Technological University (NTU), Singapore. This research is supported by DSO National Laboratories, under the project agreement No. DSOCL21238.

### v2

@@ Introduction
Hence, to improve the faithfulness of VL-CBM, we add the concept projection layer and optimize the model with our CSS objective (CSS VL-CBM). Our CSS method not only substantially increases the concept accuracy, it also closes the gap with black-box methods on image classification performance. Especially for the CUB dataset, our CSS method surpasses the black-box classifier. The “number of concept labels per class” used to train the CSS VL-CBM model for the above datasets are as CUB - 9 labels (30%), RIVAL - 8 labels (0.1%), and AwA2 - 10 labels (1.4%). The values within brackets denote the “concept labels used for training as a percentage of the total concept labels provided by the dataset”. The results show that even with such small percentages of concept labels, our CSS method can substantially increase the concept accuracy. The enhanced and faithful concept scores after improving concept alignment with CSS is presented in Figure 2 (CSS Concept Scores).
@@
WBC Attribute prediction. WBCAtt contains White Blood Cell (WBC) images annotated with 31 “cell morphology” attributes. These attributes aid in diagnosis of a variety of conditions in Hematology and Pathology. So, we directly evaluate the models on their attribute prediction accuracy and do not perform any image classification. While contemporary works (Yuksekgonul, Wang, and Zou 2023; Yang et al. 2023) use the HAM10000 (Tschandl, Rosendahl, and Kittler 2018) dataset to study VL-CBM in the medical domain, the availability of attribute labels in WBCAtt makes the latter suitable for our work. For our evaluation, we train the PLIP model (a VLM for Pathology AI), with 50% of the attribute labels. From the table, we can see that our CSS method substantially increases the attribute prediction accuracy and closes the gap with black-box attribute predictor.
Figure 3: Impact of the number of concept labels on CSS VL-CBM and LaBo.
Impact of the number of concept labels. Since our approach is semi-supervised, we further investigate the impact of the number of human-annotated concept labels on CSS VL-CBM performance, as shown in Figure 3. In this study, we also compare our method with the closest work, LaBo, where we fine-tune the final layernorm and visual projection layers of the CLIP model (which closely resembles training our concept projection layers). Both models are trained with expert concepts, and we progressively increase the number of concept labels used for training on the CUB and RIVAL datasets. From the figure, we observe that both approaches achieve similar classification accuracies. However, our CSS VL-CBM model attains relatively higher concept accuracies, especially when fewer concept labels are available. We attribute this to the use of the concept projection layer (which enhances raw CLIP scores) and the application of contrastive learning.
CBM Method Concepts Source Class Acc (%) LaBo (Yang et al. 2023) 10,000 LLM 81.9 SparseCBM (Semenov et al. 2024) 926 LLM 80.02 CDL (Zang et al. 2024) 400 LLM 83.4 LM4CV (Yan et al. 2023a) 400 LLM 81.4 CSS VL-CBM (Ours) 321 Expert 83.89
Table 3: Comparison of CSS VL-CBM with contemporary methods on CUB dataset.
Comparison with other methods
CUB. We compare our Contrastive Semi-Supervised VL-CBM method with contemporary VLM-based interpretable models on the CUB dataset, which is extensively used by the Explainable AI community to study CBMs. Following the literature, we adopt the CLIP ViT-L-14 model for this comparison. Note that this increases the number of trainable parameters in the concept projection layer. The results, presented in Table 3, show that our CSS model achieves strong classification performance despite having a smaller number of bottleneck concepts. Since other methods use LLM-generated concepts without corresponding concept labels, concept accuracy is not reported for this comparison.
CBM Method Class Acc (%) Concept Acc (%) Text2Concept 95.38 _ CSS (Ours) 98.86 72.84 PCBM 88 71.9 ECBM 91.2 85.84 CSS (Ours) 93.2 81.3
@@
Table 4: Comparison of CSS VL-CBM with other methods on RIVAL (top) and AwA2 (bottom) datasets.
@@
RIVAL and AwA2. We further compare CSS VL-CBM with contemporary methods on the animal classification datasets in Table 4. For the RIVAL dataset, CSS achieves higher classification performance compared to Text2Concept (Moayeri et al. 2023). For the AwA2 dataset, CSS achieves higher classification accuracy compared to other methods but slightly lower concept accuracy compared to ECBM. It is important to note that PCBM (Kim et al. 2023) and ECBM (Xu et al. 2023) do not use “text” concepts and optimize the vision backbone during training.
Metric CSS Scores CLIP Scores Truthfulness ($\downarrow$) 3.22 14.04 Sparseness# ($\downarrow$) 0.08 0.19 Discriminability# ($\uparrow$) 7.04 3.28
Table 5: Distributional analysis of CSS vs CLIP concept scores for CUB test. The results are averaged over instances and classes#.
Distributional Analysis of Concept Space
In addition to concept accuracy metric, we conduct a distributional analysis of the VL-CBM concept scores to explicitly demonstrate the effectiveness of our CSS method in improving concept alignment. For quantitative assessment, we extract the concept scores ($\mathbb{R}^{5790\times 312}$) from both CLIP and our CSS model for the CUB test set. We report three metrics: truthfulness: defined as the L2-distance between predicted and ground-truth concepts; sparseness: measures the intra-class standard deviation of the predicted concepts; and discriminability: defined as the L2-distance between inter-class concept score clusters. The results in Table 5 show that the proposed CSS model has improved concept alignment compared with the CLIP model. To further visually compare the difference, Figure 4 depicts a t-SNE visualization of CSS (triangles) vs. CLIP (circles) concept scores of 5 random classes on the CUB test set. Our CSS method makes the concept clusters more dense and more separable in the concept space, enhancing concept alignment by bringing them closer to the ground truth.
Figure 4: t-SNE visualization of CSS (triangles) vs CLIP (circles) concept scores of 5 random classes for CUB test.
Intervention
@@
When a CBM makes an error, one can inspect the incorrect predictions and manually adjust the classifier weights or modify the concepts to rectify these errors, thereby improving the overall accuracy. This process is known as “intervention” or “model editing.” Traditional intervention approaches are typically performed at the instance level (Oikarinen et al. 2023; Yuksekgonul, Wang, and Zou 2023; Koh et al. 2020), where a few erroneous images are arbitrarily chosen for intervention. However, for fine-grained multi-class classification problems like CUB, which have many classes (some bird species are visually similar) and relatively few samples per class, selection of error images become non-trivial. Therefore, we propose a class-level intervention procedure where we first identify the confounding classes and then choose the error images from these classes for intervention. Confounding classes are bird species that are visually similar yet semantically different, and the model makes the most errors for these challenging classes. This class-level intervention is carried out in the following steps.
Step 1: Identify the confounding classes. We compute a $K\times K$ error matrix by evaluating the trained CSS VL-CBM model on the test set, where $K$ is the number of classes. The error matrix provides the error distribution for each class, and the matrix element with the highest value indicates the confounding class pairs. In our experiments, we identified two pairs of confounding classes: California Gull - Western Gull and Common Tern - Arctic Tern. The total number of classification errors for these confounding classes before intervention is presented in Table 6.
Step 2: Inspecting the concept space. Next, we inspect the concept scores of the error images in these confounding class pairs. We find that the concept score distribution of the confounding classes is nearly identical, making it difficult to discriminate between them. Even the averaged ground-truth concept labels for the confounding classes are highly similar. Our model suffers from Type-2 error (Oikarinen et al. 2023), meaning the number of bottleneck concepts and their score distributions are insufficient to effectively discriminate the samples.
Before Interv.   After Interv. Total error for CG (30) 27 14 Total error for WG (30) 15 9 CG misclassified as WG 13 4 WG misclassified as CG 3 1 Total error for CT (30) 21 11 Total error for AT (29) 3 2 CT misclassified as AT 12 6 AT misclassified as CT 0 0 Classification Accuracy (%) 81.45 82.57
@@
Step 3: Expanding the concept set. To mitigate the Type-2 error and to improve the overall performance, we expand the concept set by leveraging the concepts provided by LaBo (Yang et al. 2023). Note that the new concepts to be added should be visually groundable by the CLIP model and hence we use the “filtered” concept set of LaBo. We choose $top-32$ concepts per confounding class and expand the concept set from the original 312 concepts to 440 concepts. Formally, let $T^{\prime}$ be the new set of 128 text concepts and $C^{\prime}_{i}=E_{I}(x_{i})\cdot E_{T}(T^{\prime})^{\intercal}$ be the CLIP concept scores for it. Now the concept scores for all the 440 concepts for an image $x_{i}$ is given by $C^{\prime}=$ Concat($C_{i},C^{\prime}_{i}$). During this intervention procedure we do not train a new concept projection layer for the expanded concept set.
@@
Limitations
Despite the numerous advantages, VL-CBMs have a few potential limitations.
Ineffable concepts. While natural language offers a convenient way to build interpretable models with high-level abstract concepts, its expressive power can be limited. The subtle visual cues required for certain classification tasks, such as face recognition, might be difficult to articulate in words.
Unknown concepts. VL-CBMs assume that the complete set of salient concepts required for solving the classification task is known before training. However, this assumption is not always feasible. In such cases, models that can “discover” unknown concepts (Shang et al. 2024) during the course of training are more suitable.
Locality Faithfulness. Recent works have observed that even when CBMs predict the concept correctly, they often fail to ‘respect localities’ (Raman et al. 2024). This suggests that the concept predictor might rely on spurious correlations instead of relevant visual/spatial features. Such a study on locality faithfulness in VL-CBMs is largely unexplored.
Conclusion
In this paper, we investigated the faithfulness of VL-CBM for expert concepts and found that VLMs like CLIP have poor concept alignment. Hence, we proposed a novel Contrastive Semi-Supervised approach that improved concept alignment with fewer human-annotated concept labels. Our work will primarily aid the Explainable AI community in building interpretable models that align with human preferences (concepts) yet require minimal supervision. We also introduced a class-level intervention procedure for fine-grain classification problems that reduced the error for the confounding classes to increase the overall performance of the model.